OpenStudy (anonymous):

Solve A = πr + 4t for t.

OpenStudy (anonymous):

\[t = \frac{A - \pi r}{4}\]

OpenStudy (anonymous):

Thank Yu So Much (:

OpenStudy (anonymous):

Solve 3x – 5y = 15 for y.